How do you convert 7534 base 8 to base 10?
Found 2 solutions by dabanfield, nerdybill:
Answer by dabanfield(803)

How do you convert 7534 base 8 to base 10?
7534 (base 8) =
4*8^0 + 3*8^1 + 5*8^2 + 7*8^3 =
4 + 3*8 + 5*64 + 7*512 =
4 + 24 + 320 + 3584 = 3932 (base 10)

Answer by nerdybill(6958)

How do you convert 7534 base 8 to base 10?
.
It would be:
7*8^3 + 5*8^2 + 3*8^1 + 4*8^0
= 7*512 + 5*64 + 3*8 + 4*1
= 3584 + 320 + 24 + 4
= 3932